# Semantics-Aligned, Curriculum-Driven, and Reasoning-Enhanced Vulnerability Repair Framework

### v1

@@
Abstract.
Current learning-based Automated Vulnerability Repair (AVR) approaches, while promising, often fail to generalize effectively in real-world scenarios. Our diagnostic analysis reveals three fundamental weaknesses in state-of-the-art AVR approaches: (1) limited cross-repository generalization, with performance drops on unseen codebases; (2) inability to capture long-range dependencies, causing a performance degradation on complex, multi-hunk repairs; and (3) over-reliance on superficial lexical patterns, leading to significant performance drops on vulnerabilities with minor syntactic variations like variable renaming.
To address these limitations, we propose SeCuRepair, a semantics-aligned, curriculum-driven, and reasoning-enhanced framework for vulnerability repair. At its core, SeCuRepair adopts a reason-then-edit paradigm, requiring the model to articulate why and how a vulnerability should be fixed before generating the patch. This explicit reasoning enforces a genuine understanding of repair logic rather than superficial memorization of lexical patterns. SeCuRepair also moves beyond traditional supervised fine-tuning and employs semantics-aware reinforcement learning, rewarding patches for their syntactic and semantic alignment with the oracle patch rather than mere token overlap. Complementing this, a difficulty-aware curriculum progressively trains the model, starting with simple fixes and advancing to complex, multi-hunk coordinated edits.
We evaluate SeCuRepair on strict, repository-level splits of BigVul and newly crafted PrimeVulAVR datasets. SeCuRepair significantly outperforms all baselines, surpassing the best-performing baselines by 34.52% on BigVul and 31.52% on PrimeVulAVR in terms of CodeBLEU, respectively. Our human evaluation of patch correctness further shows that SeCuRepair generates 16% more workable patches than the best-performing baseline, GPT-4o. Comprehensive ablation studies further confirm that each component of our framework contributes to its final performance.
Vulnerability Repair, Large Language Model, Reinforcement Learning
copyright: acmlicensedjournalyear: 2025ccs: Security and privacy Software and application security
1. Introduction
Software vulnerabilities are weaknesses in software that attackers can exploit to violate security policies, e.g., confidentiality, integrity, or availability (nist7435, ). The number of reported software vulnerabilities continues to rise steadily. In 2024, the National Vulnerability Database (NVD) published 40,009 Common Vulnerabilities and Exposures (CVE) records, representing a 38% increase compared to 2023. Despite this growing trend in reporting, the verification and remediation of CVEs remain challenging and have not kept pace with this growth. There are still more than 25,900 CVEs awaiting initial analysis by NVD (nvd_dashboard, ) (as of 01/09/2025) and they often take weeks to be fixed (cve_fix_time1, ; cve_fix_time2, ), highlighting systemic triage pressure and slower downstream remediation. As a result, these trends motivate automatic vulnerability repair approaches to help in vulnerability patching.
A number of state-of-the-art and learning-based approaches have been proposed for automatic vulnerability repair (AVR), which frame AVR as function-level sequence-to-sequence generation (fu2022vulrepair, ). Given a vulnerable function and optional auxiliary contexts, an AVR model is trained to generate a patch. Representative AVR approaches, including VREPAIR (vrepair, ), VulMaster (vulmaster, ), and FAVOR (favor, ), perform supervised fine-tuning (SFT) with single-reference patches, enriched with auxiliary signals. In particular, VREPAIR (vrepair, ) introduces a transfer-learning pipeline that pretrains on large bug-fixing corpora. Also, VulMaster (vulmaster, ) accommodates long function input and integrates CWE knowledge to guide repair. Moreover, FAVOR (favor, ) augments the input with historical patch patterns.
Despite their success, we identify three key limitations of existing learning-based AVR approches.
Figure 1. Limitation 1: Performance comparison of VulMaster, FAVOR, and CodeT5 under random split and repository-level split. The performance of selected models degrades substantially under repository-level split by up to 91.6% in terms of Exact Match and up to 29.7% in terms of CodeBLEU.
((a)) Perfect Repair Patches Generated by VulMaster Fine-tuned on Random-split Dataset.
((b)) Incorrect Patches Generated by VulMaster after Renaming One Local Variable plenbytes to nbytes.
Figure 2. VulMaster fails to generate the buffer over-read guard when a local variable is renamed with a similar naming convention.
Limitation 1: Limited cross-repository generalization. Prevailing AVR approaches (vrepair, ; vulmaster, ; favor, ) are evaluated on random-split datasets, which overlook cross-repository generalization. Given the scale of real-world code (over 420M GitHub repositories (githubstatistics, )) versus the limited available vulnerability corpora, a practical AVR approach must generalize to unseen projects. To assess this, we re-evaluate existing AVR approaches by splitting the dataset without repository overlap between the training and test sets (details in §2.2.1). As shown in Fig.1, with this realistic setup, we reveal a substantial generalization gap of existing learning-based AVR approaches: they can not maintain high effectiveness as before, with CodeBLEU dropping by up to 29.7% and Exact-Match up to 91.6%.
Additionally, a qualitative case study on CVE-2018-16228 (Fig. 2) further highlights the brittleness of prevailing AVR approaches on semantically equivalent vulnerable code. While VulMaster correctly patches the original function, it fails after syntactic-modified but semantics-preserving edits (e.g., identifier renaming), revealing a critical sensitivity to syntactic variations.
We attribute this poor cross-repository generalization and syntactic brittleness to the models’ overfitting on spurious and repository-specific lexical correlations rather than learning syntactic and semantic repair patterns. The prevalent training objective in AVR approaches, maximizing token likelihood against a single oracle patch, provides no supervision to distinguish meaningful semantic edits from superficial lexical cues (e.g., specific variable names or API idioms). The loss function simply rewards matching the oracle’s tokens, allowing lexical patterns to be learned instead of the underlying repair logic. These findings motivate the advancement of AVR approaches that jointly account for syntactic and semantic features in repair, while emphasizing the necessity of rigorous, cross-repository evaluation.
Figure 3. Limitation 2: The performance of existing AVR approaches drops as the number of hunks increases.
Limitation 2: Unsolved multi-hunk repair optimization. Many real-world vulnerabilities require multi-hunk repairs: a set of coordinated edits across different, often distant, hunks of a function. For example, in vulnerability datasets like BigVul and PrimeVul, 39.33% and 47.07% of the vulnerable functions require multi-hunk repairs, respectively. Unfortunately, as shown in Fig. 3 (detailed in Sec 2.2.3), the performance of representative AVR approaches generally degrades with the number of hunks. Specifically, VulMaster, FAVOR, and CodeT5 performance in terms of CodeBLEU drops by 19.75%, 11.27%, and 13.22%, respectively, when moving from 1 to 2 hunks. Our finding highlights the need for AVR approaches to perform better for multi-hunk repairs, which are common in real-world scenarios.
Limitation 3: Explicit repair reasoning, underexplored yet broadly useful. Current AVR approaches operate like black-box translators: converting vulnerable code directly into a patched version without generating any explicit reasoning about the fix. They do not articulate the vulnerability’s root cause, the required logical changes, or a step-by-step repair plan. They force the models to generate a patch based only on local patterns learned from the data, without a mechanism to enforce the global consistency of the fix.
Recent work suggests a path forward: structured reasoning approaches have improved performance in vulnerability detection (weyssow2025r2vul0, ), while causal learning methods can eliminate reliance on spurious features (rahman2024causal, ). We argue that AVR can benefit by making repair intent explicit. Rather than directly generating patches, models should first articulate a repair plan, then implement it. Yet this reasoning-then-edit approach remains largely unexplored for AVR.
Our Work. We present SeCuRepair, a Semantic and syntactic-aligned, Curriculum-driven, Reasoning-enhanced AVR framework, to address the three limitations above directly: 1) Rewarding semantically and syntactically aligned patches (addressing Limitation 1). To combat overfitting on spurious token patterns, SeCuRepair’s training objective moves beyond SFT’s token-matching. Rather, it uses a syntactic- and semantic-aligned reinforcement learning (RL) objective that rewards generated patches for their syntactic (Abstract Syntax Tree) and semantic (Data Flow Graph) similarity to the oracle patch. This encourages the model to learn the behavioral repair logic rather than just mimicking lexical patterns. 2) Curriculum scheduling (addressing Limitation 2). To master multi-hunk repairs, SeCuRepair employs a difficulty-aware curriculum. The model is trained progressively, starting with patches with less number of hunks before advancing to patches requiring multiple, non-contiguous edits. This approach allows the model to master foundational repair patterns before tackling the intricate dependencies of more complex vulnerabilities. 3) Reason-then-edit generation (addressing Limitation 3). To prevent inconsistent or partial fixes, SeCuRepair adopts a reason–then–edit paradigm. It first generates a concise, natural-language plan for the repair and then produces the code conditioned on that plan. This provides a strong guide for generating globally consistent and logically sound edits.
We conduct a comprehensive evaluation of SeCuRepair on two benchmarks: the widely-used BigVul dataset (fan2020ac, ) and PrimeVulAVR, a dataset we adapt from vulnerability detection and extend for repair. We compare SeCuRepair against state-of-the-art AVR approaches, including VulMaster (vulmaster, ), FAVOR (favor, ), and the general-purpose GPT-4o. The results demonstrate the superiority of our approach. SeCuRepair outperforms the strongest baseline GPT-4o by a significant margin, achieving relative CodeBLEU improvements of 34.52% on BigVul and 31.52% on PrimeVulAVR. Meanwhile, our human evaluation of patch correctness shows that SeCuRepair generates 16% more workable patches than the best-performing baseline, GPT-4o.
Furthermore, a series of ablation studies confirms the individual effectiveness of each core component within the SeCuRepair framework. In summary, the main contributions are as follows:
  • Empirical Analysis. We are the first to systematically analyze the challenges of realistic cross-repository generalization and multi-hunk repairs in existing AVR approaches.
  • Approach. We design and implement SeCuRepair, a novel AVR framework that integrates syntactic- and semantic-aligned reinforcement learning, curriculum-based training, and a reason-then-edit workflow to address these identified challenges.
  • Evaluation. Our extensive experiments demonstrate that SeCuRepair consistently outperforms existing methods on BigVul and PrimeVulAVR.
2. Problem Statement and Empirical Analysis
In this section, we first formally define the problem of AVR and then empirically analyze the challenges in existing AVR approaches: cross-repository generalization, multi-hunk repair performance, and brittleness to semantics-preserving code refactoring.
2.1. Problem Definition
Following prior works (vrepair, ; vulmaster, ; favor, ), we formulate AVR as a function-level sequence-to-sequence generation task. Specifically, given a vulnerable function $X$ and auxiliary information $L$ (e.g., CWE information and localization information), the goal is to produce a repair function $Y$ that fixes the vulnerability in $X$. An AVR model is trained on a dataset $\mathcal{D}=\{(X_{i},L_{i},Y_{i})\}_{i=1}^{N}$ of $N$ examples. Existing AVR approaches (vrepair, ; vulmaster, ; favor, ) typically optimize the model with supervised-fine-tuning (SFT) to maximize the likelihood of the reference patch $Y_{i}$ given $X_{i}$ and $L_{i}$:
(1) $\mathcal{L}_{SFT}=-\sum_{i=1}^{N}\log P(Y_{i}|X_{i},L_{i};\theta)$
where $\theta$ denotes the model parameters.
2.2. Empirical Analysis
To systematically assess the limitations of current AVR approaches, we conduct a series of controlled experiments using a consistent evaluation protocol. We focus on three representative SFT-based baselines, i.e., VulMaster (vulmaster, ), FAVOR (favor, ), and their base LLM CodeT5 (wang2021codet5, ). Following standard practice (vulmaster, ; favor, ; vrepair, ), we report performance using CodeBLEU (codebleu, ) and Exact Match (EM) scores. In addition, we also conduct manual evaluation to assess the correctness of generated patches and the quality of the reasoning. Each subsequent subsection introduces one experimental variant targeting (i) cross-repository generalization, (ii) multi-hunk repair, and (iii) robustness to semantics-preserving refactoring.
2.2.1. Cross-Repository Generalization Gap
Setup. As noted in Section 1, AVR models must generalize beyond the repositories seen in training. We formalize this through a cross-repository evaluation protocol. We evaluate the baselines on the common dataset used by VulMaster and FAVOR’s dataset: BigVul dataset (fan2020ac, ). We compare their performance under two distinct data splitting strategies: the conventional random split (using the same split as FAVOR) and our strict repository-level split. In the repository-level split, all functions from a given project are confined to a single set (training, validation, or test). To ensure a fair comparison, we fine-tune all models for each strategy. To prevent data contamination, we exclude the pre-trained adaptor modules of VulMaster, as their training relies on a bug-fixing corpus with known project-level overlaps with the BigVul test set.
The performance of selected AVR approaches drops substantially under repository-level split by up to 89.86% in terms of Exact Match and up to 29.70% in terms of CodeBLEU.
2.2.2. Brittleness to Semantics-Preserving Code Refactoring
Setup. To probe whether AVR approaches learn real semantic repair patterns or merely overfit to lexical cues, we test their robustness under a simple refactoring: renaming local variables while preserving semantics. We randomly selected 10 cases where baseline models (trained and evaluated under random splits) correctly generated a patch, then applied consistent variable renamings.
Observation. All studied models are highly sensitive to these minor changes, failing to generate functionally equivalent patches in 40–70% of cases, based on manual inspection by the first author. Figure 2 provides an example of this failure. In the original patch (a), VulMaster correctly inserts a safety guard to prevent a buffer over-read. However, after simply renaming the variable plenbytes to nbytes (b), the model fails. It does not generate the equivalent safety check but instead hallucinates an irrelevant function call, leaving the vulnerability unresolved.
This case study provides evidence that the model has not learned the underlying semantic rule, i.e., “the return value of the decoding function must be checked before proceeding.” Instead, it learns a superficial correlation between the specific variable name plenbytes and the patch. While in limited scale, the consistent failures strongly motivates the need for an AVR learning framework that prioritizes syntactic and semantic understanding of repair patterns over simple lexical mimicry.
The Brittleness of the Exact Match Metric. The case study above also highlights the fundamental brittleness of the Exact Match (EM) metric for evaluating AVR. EM is fundamentally sensitive to syntactic variations that preserve the code’s semantics. For example, a logically equivalent change like rewriting if (a > b) to if (b < a) would still result in an EM score of 0. Indeed, an ideal AVR approach capable of semantic repairing should be able to generate a diverse set of semantically correct and equivalent patches, most of which would be wrongly evaluated by EM as they are lexically different from the single oracle patch. This strictness makes EM an unreliable indicator of a patch’s correctness and ill-suited for evaluating advanced, semantics-aware AVR approaches. In contrast, CodeBLEU considers AST similarity, making it robust to syntactic variations. Therefore, we consider CodeBLEU a more meaningful metric for AVR and use it as our primary measure in the following sections.
The failure of existing AVR approaches on simple, semantics-preserving changes reveals their over-reliance on lexical patterns over semantic repair patterns. This same flaw suggests EM may have limitations as a metric for AVR, as it is confined to judging lexical equivalence with the oracle patch.
2.2.3. Suboptimal on Multi-Hunk Repair
Setup. Real-world vulnerability fixes frequently span multiple, non-contiguous regions (hunks) that must be edited consistently. Prior work FAVOR (favor, )’s case study notes that multi-hunk repair significantly increases AVR complexity and is a common failure mode for sequence-to-sequence models. To quantify this, we evaluate baseline models on our repository-split test set, stratifying the results by the number of hunks in each function.
The performance of AVR models decreases as the complexity of repairs increases. All studied models become progressively less effective as the number of required edits (hunks) increases.
3. Approach
Inspired by the analysis in Section 2.2, we propose SeCuRepair, a framework designed to learn the underlying syntactic and semantic logic of vulnerability repairs rather than imitating superficial, repository-specific lexical patterns. The design of SeCuRepair is guided by three core principles:
  • A Reason-then-Edit Paradigm: To ensure logical consistency in repairs, the model is trained to first generate an explicit, natural-language plan that reasons the vulnerability and the proposed fix before it writes the code.
  • Semantics-Aware Optimization: To promote robust generalization, the model is optimized using rewards that measure the syntactic and semantic correctness of a patch, moving beyond simple textual similarity to a single reference solution.
  • Difficulty-based Curriculum Learning: To master complex, multi-hunk fixes, the model is trained on a curriculum of increasing difficulty, allowing it to learn foundational repair patterns before advancing to more challenging scenarios.

### v2

# SeCuRepair: Semantics-Aligned, Curriculum-Driven, and Reasoning-Enhanced Vulnerability Repair Framework

@@
Abstract
The rapid accumulation of software vulnerabilities has outpaced manual remediation, creating an urgent need for Automated Vulnerability Repair (AVR). However, existing methods suffer from syntactic overfitting, mimicking surface forms without understanding the underlying repair logic, and fail to generalize to complex fixes. To transcend these limitations, we propose SeCuRepair, a reliable, scalable, and efficient RL-based AVR framework. By introducing a semantic-aware reward, SeCuRepair optimizes for code semantic equivalence rather than lexical mimicry. Furthermore, SeCuRepair incorporates an expert-aligned reasoning mechanism that explicitly grounds patch generation in a structured diagnosis. Finally, SeCuRepair introduces a difficulty-based curriculum that progressively disentangles the optimization barriers of entangled multi-hunk repairs. Extensive evaluations on a rigorous repository-level split demonstrate that SeCuRepair significantly outperforms state-of-the-art baselines via automatic evaluation and human study.
SeCuRepair: Semantics-Aligned, Curriculum-Driven, and Reasoning-Enhanced Vulnerability Repair Framework
Chengran Yang1   Ting Zhang2   Jinfeng Jiang1   Xin Zhou1   Haoye Tian3   Mingzhe Du4 Jieke Shi1Junkai Chen1Yikun Li1Eng Lieh Ouh1Lwin Khin Shar1David Lo1 1Singapore Management University, Singapore  2Monash University, Australia 3Aalto University, Finland  4National University of Singapore, Singapore
1 Introduction
The rapid accumulation of software vulnerabilities has overwhelmed manual remediation capacities. The National Vulnerability Database (NVD) published 49,230 Common Vulnerabilities and Exposures (CVE) records in 2025 alone, yet due to slow human remediation, an overall total of 27,900 CVEs remain unanalyzed 20. With manual fixes often requiring weeks  Forsgren et al. (2021), there is a critical need for effective Automatic Vulnerability Repair (AVR).
To address this backlog, existing approaches rely on Supervised Fine-Tuning (SFT) Chen et al. (2022); Zhou et al. (2024); Dong et al. (2025), training models to map vulnerable code to a specific human-written patch. However, this formulation is fundamentally misaligned with the intrinsic flexibility of AVR: real-world vulnerability admits diverse semantically equivalent fixes, while SFT optimizes for the exact reconstruction of a unique syntactic solution. This rigid objective penalizes generating semantically equivalent but syntactically different patches, forcing models to overfit to superficial coding patterns.
Reinforcement learning (RL) transcends SFT’s rigidity by shifting the optimization objective from token-level imitation to outcome-driven exploration Shao et al. (2024). However, applying RL to AVR is impeded by two fundamental challenges: the lack of a reliable reward signal and the complexity of the exploration landscape.
First, constructing a reliable reward signal for AVR is difficult. While execution-based rewards are standard in Code LLMs Guo et al. (2025); Hui et al. (2024), real-world AVR datasets lack the large-scale build environments and test cases required for dynamic verification. Consequently, execution-free alternatives prove inadequate in AVR: 1) Lexical Rewards (e.g., SWE-RL Wei et al. (2025)) rely on string similarity to the reference, which effectively reintroduces SFT’s rigidity by penalizing valid but syntactically different patches; 2) LLM-as-Judge is unreliable in AVR, often failing to distinguish between vulnerable and patched functions Ding et al. (2024), thereby providing noisy and indiscriminative signals. This leaves a critical gap for a reward design that is execution-free yet rewards semantic equivalence.
Second, efficient exploration is hindered by the inherent complexity of vulnerability repair. Direct RL on AVR often yields sub-optimal policies due to two critical characteristics: 1) Reasoning-dependent Nature: effective repair is an inherent chain-of-thought process, requiring the inference of underlying vulnerability logic and analyzing of the root cause before generating a patch. Bypassing this reasoning process leads to a reasoning gap, rendering the exploration phase unstable. 2) Entangled Learning Objectives: complex repairs spanning multiple code hunks (i.e., non-contiguous edited regions in a function) require the model to simultaneously master precise local edits (for security) and global structural synchronization (for syntactic validity). Attempting to optimize these entangled objectives concurrently overwhelms the exploration process, often trapping the LLM exploration in local optima.
To bridge these gaps, we propose SeCuRepair, which enables reliable, scalable, and efficient RL-based vulnerability repair. To bridge the gap of reliable reward, we propose a multi-grained reward. We proxy patch quality through a composite metric of lexical (BLEU), syntactic (AST), and semantic (Data and Control Flow Graph) similarity to human reference. This formulation anchors the reward in semantic invariants, rewarding semantically equivalent but syntactically diverse patches. Meanwhile, it enables scalable training across large-scale, real-world AVR datasets where build environments are missing.
For efficient exploration, we propose a reasoning-enhanced initialization followed by a curriculum training strategy. Specifically, we employ an expert-aligned reasoning protocol to initialize the policy, ensuring LLM executes a diagnostic process before attempt to fix. Meanwhile, we organize training samples by complexity, utilizing the number of code hunks as a proxy. This curriculum guides the LLM to first master isolated local fixes by starting with single-hunk samples, before progressively advancing to tackle global structural coordination in complex scenarios.
We evaluate the efficacy of SeCuRepair on established benchmark BigVul Fan et al. (2020) and our newly proposed PrimeVulAVR, which employing a strict repository-level splitting strategy to avoid data leakage. SeCuRepair demonstrates significant superiority over state-of-the-art baselines, achieving a 37.21% improvement in CodeBLEU on BigVul and 33.58% on PrimeVulAVR. Human evaluation further confirms SeCuRepair performs on par with GPT-4o.
In summary, our main contributions are as follows: 1) We propose SeCuRepair, an RL-based framework that transcends SFT rigidity via a semantic-aware reward. 2) We devise a curriculum training strategy combined with reasoning-enhanced distillation for efficient exploration. 3) We establish a rigorous cross-repository evaluation protocol to eliminate the data leakage inherent in random splits, and introduce a new AVR benchmark PrimeVulAVR.
2 Related Work
AVR approaches. Rule-based methods, ranging from template-guided Lin et al. (2007) to constraint-based solvers Zhang et al. (2022); Gao et al. (2021). They rely on either rigid, manually defined patterns or expensive runtime execution, making them ineffective for the majority of vulnerabilities where build environments are unavailable. Leading learning-based approaches such as VulMaster Zhou et al. (2024) and FAVOR Dong et al. (2025) improve repair performance by incorporating external knowledge and structural context based on the SFT objective. Yet, their reliance on SFT inherently encourages lexical mimicry over semantic understanding.
RL in software engineering. Reinforcement Learning has shown promise in various software engineering tasks Guo et al. (2025); Hui et al. (2024); Wei et al. (2025); Yang et al. (2024), typically relying on execution feedback as reward. However, this is impractical for AVR due to the scarcity of large-scale build environments. Existing execution-free alternatives also fall short. Lexical rewards (e.g., SWE-RL Wei et al. (2025)) measure string similarity to a reference, but this rigid metric penalizes syntactically diverse yet semantically equivalent solutions, providing a biased signal. Model-based rewards Dutta et al. (2024); Sghaier et al. (2025) employ LLMs as judges, but these prove unreliable for security tasks, as even SOTA models struggle to distinguish between vulnerable code and its patched version Ding et al. (2024), which is the core of AVR rewarding. In contrast, our approach bridges this gap by introducing a multi-grained, semantic-aware reward, proxying patch quality through static analysis of syntactic and semantic equivalence.
3 Problem Statement and Motivation
Figure 1: Syntactic overfitting of SFT: The SFT model generates a correct fix for the buffer over-read guard with the original input, but fails to generalize when a local variable is simply renamed from plenbytes to nbytes.
In this section, we formulate the task of AVR and analyze the limitations of existing AVR approaches: 1) Syntactic Overfitting; 2) Suboptimization on Complex Repair; and 3) Biased Evaluation Setting.
3.1 Problem Definition
Following prior works Chen et al. (2022); Zhou et al. (2024); Dong et al. (2025), we formulate AVR as a function-level sequence-to-sequence generation task. Given a vulnerable function $X$ and auxiliary information $L$ (e.g., information of vulnerability localization and Common Weakness Enumeration), the goal is to generate a patched function $Y$ that fixes the vulnerability in $X$.
3.2 Syntactic Overfitting of SFT in AVR
To probe whether existing SFT approaches Wang et al. (2021); Zhou et al. (2024); Dong et al. (2025) learn semantic repair patterns or overfit to syntactic cues, we follow Rahman et al. (2024) and test their performance under a simple refactoring: renaming local variables of input while preserving code semantics (see Appendix B for details). We observe that all SFT models are sensitive to these minor changes, failing to generate valid patches in 40–70% of cases. Figure 1 provides an example of this failure. A simple variable renaming (plenbytes $\to$ nbytes) causes VulMaster to regress from a correct fix to an irrelevant hallucination. This provides empirical evidence that SFT models heavily rely on superficial variable correlations. In this paper, we propose an RL training framework that rewards semantically equivalent patches to mitigate syntactic overfitting.
3.3 Suboptimization on Multi-hunk Repair
A critical oversight in current AVR research is the lack of optimization for complex, multi-hunk repair. Real-world fixes often require changes across multiple discontinuous code regions, yet existing methods treat all samples uniformly. To quantify this impact, we stratified the performance of representative baselines, VulMaster Zhou et al. (2024) and FAVOR Dong et al. (2025), by the number of repair hunks in human reference. Performance drops with complexity: even moving from 1 to 2 hunks causes 11% to 20% CodeBLEU drops across all approaches (see Figure 6 for full results).
Our manual inspection reveals that 55% of multi-hunk failures stem from global inconsistencies. We attribute this failure to entangled learning objectives: the model faces the dual burden of executing precise local security edits while simultaneously maintaining global syntactic synchronization. We propose a curriculum-training strategy to mitigate this issue, which guides the LLM to first master local fixes by starting with single-hunk samples. This setting smooths the optimization landscape, allowing the model to learn patch patterns without global interference Bengio et al. (2009). We then introduce multi-hunk samples, allowing LLM to focus more on aligning global dependencies.
3.4 Data Leakage in Conventional Evaluation
Existing AVR approaches typically rely on function-level random splitting Zhou et al. (2024); Dong et al. (2025). However, this setting poses a significant data leakage risk Zhang et al. (2025), as it allows models to exploit project-specific patterns or "peek" at future fixes within the same repository. We quantify this threat by comparing model performance under the random-split versus a strict cross-repository protocol, where all functions from a specific project are confined to a single split (train/val/test). The results reveal a severe generalization gap. Under the strict repository-level split, performance degrades sharply across all baselines: CodeBLEU scores drop by a relative 21.49%–29.70%, while exact match scores plummet by 87.88% to 89.86% (see Appendix D for full details). These findings highlight the leakage issue in a random-split setting, inspiring us to use a cross-repository setting in the following sections.
4 Approach
Driven by the above limitations, we propose SeCuRepair, an RL-based AVR framework to enforce LLM to learn the underlying logic of vulnerability repair. The design of SeCuRepair is guided by three core principles:
1) Reasoning-enhanced Initialization: Vulnerability repair inherently demands complex logic: root cause analysis, vulnerability type classification, and devising a fix strategy. To mitigate the optimization instability caused by this reasoning gap, we warm-start the model via knowledge distillation, leveraging a structured protocol that simulates a security-expert cognitive workflow.
2) Semantics-Aware Optimization: To overcome the syntactic overfitting of SFT, we propose a multi-grained static reward signal for patch quality, which enables the measurement of semantic equivalence.
3) Difficulty-based Curriculum Learning : To dismantle the optimization barrier in complex multi-hunk repairs, we employ a difficulty-based curriculum. This approach decouples the learning process by progressively increasing task complexity, allowing the model to master local security edits before tackling global synchronization.